OpenStudy (anonymous):

Help Please! (: Solve each equation by completing the square:

OpenStudy (anonymous):

\[k^2-8k-48=0\]

OpenStudy (compassionate):

Why not use the quadratic formula? Or - you can factor the equation. Can you do that for me?

OpenStudy (dumbcow):

oh come on, completing the square is so much fun though :)

OpenStudy (anonymous):

Haha! @dumbcow :P

OpenStudy (compassionate):

wut. No, it's a trinomial. You can factor it into two binomials.

OpenStudy (compassionate):

Factor. Have you taken 8th grade math, hun? Do you know how to factor a trinomial?

OpenStudy (jim766):

if her directions say to complete the square, that is the way she should do it...

OpenStudy (dumbcow):

\[k^{2} -8k = 48\] \[k^{2} - 8k+(\frac{8}{2})^{2} = 48 + (\frac{8}{2})^{2}\] \[k^{2} -8k +16 = 64\] \[(k-4)^{2} = 64\] \[k-4 = \pm 8\] \[k = -4 , 12\]

OpenStudy (anonymous):

@dumbcow Thanks so much! Much appreciated!! GBU!! (:

OpenStudy (dumbcow):

hopefully you understand the steps and @Compassionate has a point , at this stage of math you should be able to easily factor this quadratic as well

OpenStudy (compassionate):

Exactly. Like Dumbcow said. I was surprised that you're working on Quadratics and don't even know how to factor a trinomial.

OpenStudy (dumbcow):

\[k^{2} -8k -48 \] 2 num that multiply to -48 but add to -8 -12*4 = -48 -12 +4 = -8 \[\rightarrow (k-12)(k+4)\]